# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

### v1

@@ Abstract
With the rapid growth of multi-modal data from social media, short video platforms, and e-commerce, content-based retrieval has become essential for efficiently searching and utilizing heterogeneous information. Over time, retrieval techniques have evolved from Unimodal Retrieval (UR) to Cross-modal Retrieval (CR) and, more recently, to Composed Multi-modal Retrieval (CMR). CMR enables users to retrieve images or videos by integrating a reference visual input with textual modifications, enhancing search flexibility and precision. This paper provides a comprehensive review of CMR, covering its fundamental challenges, technical advancements, and categorization into supervised, zero-shot, and semi-supervised learning paradigms. We discuss key research directions, including data augmentation, model architecture, and loss optimization in supervised CMR, as well as transformation frameworks and external knowledge integration in zero-shot CMR. Additionally, we highlight the application potential of CMR in composed image retrieval, video retrieval, and person retrieval, which have significant implications for e-commerce, online search, and public security. Given its ability to refine and personalize search experiences, CMR is poised to become a pivotal technology in next-generation retrieval systems. A curated list of related works and resources is available at: Awesome Composed Multi-modal Retrieval.
@@
Composed Multi-modal Retrieval (CMR), Cross-modal Retrieval, Multi-modal Learning, and Vision-Language Models
@@ 1 Introduction
Nowadays, the unprecedented prosperity of social media, short video platforms, and e-commerce (such as TikTok, Facebook, Weibo, YouTube, Amazon, and Alibaba) has greatly promoted the rapid growth of multi-modal data, encompassing various modalities such as texts, images, and videos. Faced with this data flood composed of heterogeneous information, content-based retrieval, as an information processing technology to search and utilize these vast resources, not only plays a core role in the field of e-commerce, such as achieving accurate matching of goods, but also greatly enriches the user experience of the social media platform, allowing users to quickly and accurately find the content of interest. Due to its widespread applications in everyday life that almost impact everyone, content-based retrieval has attracted great attention from both academia and industry, whose research significance is self-evident.
@@
Generally, the evolution of content-based retrieval technology has witnessed the transformation from Unimodal Retrieval (UR) to Cross-modal Retrieval (CR), and then to Composed Multi-modal Retrieval (CMR). Compared with the early-stage unimodal retrieval, which was limited to querying information within the same modality, e.g., image search for images, cross-modal retrieval has achieved remarkable accuracy and widespread application in the present era. This enables the search for semantically relevant content in one modality based on the instance query from another modality, e.g., using text search images or videos, allowing users to make full use of these heterogeneous data. In recent years, composed multi-modal retrieval has emerged as a thriving content-based retrieval technology. Within this technical framework, the system aims to discover images/videos that not only bear resemblance to the given reference image/video but also allow for specific modifications based on the provided textual feedback from the user. In this sense, CMR pioneers an advanced level of interactive and conditional retrieval mechanism, leveraging deep integration of visual and linguistic information. This integration greatly enhances the flexibility and precision of user-expressed search intents, injecting new vitality into domains such as internet search and e-commerce. Consequently, CMR exhibits vast potential and far-reaching impact as the next-generation content-based retrieval engine in real-world application scenarios.
@@
A core of CMR is that it requires the synergistic understanding and composing of both input vision and language information as the multi-modal query. The earliest closely related studies of CMR are in the field of attribute-based fashion image retrieval, where the key difference is that the textual feedback in attribute-based fashion image retrieval is limited to a single attribute (e.g., ’mini’, ’white’, ’red’), while CMR is the natural language with multiple words (e.g., ’showing this animal of the input image facing the camera under sunlight’), which is more flexible yet challenging compared to the predefined set of attribute values. The pioneering CMR works are proposed in, where the input query is specified in the form of an image plus some natural language that describes desired modifications to the input image, and led a series of subsequent approaches. Current research in CMR is primarily focused on three main patterns: (1) supervised learning-based CMR (SL-CMR), which focuses on how to design a better combination mechanism of vision and language through supervised training of annotated data, (2) zero-shot learning-based CMR (ZSL-CMR), which focuses on how to simulate and build a visual-linguistic multi-modal information combination framework without annotated data, and (3) semi-supervised learning-based CMR (SSL-CMR), which focuses on how to enhance the learning of visual-linguistic combination via automatic generated data under a semi-supervised manner.
@@
For the combined multi-modal retrieval models based on supervised learning, a notable characteristic is the requirement of annotated triplet data $(I^{r},T^{r},I^{t})$, where $I^{r}$ denotes the reference query image, $T^{r}$ denotes the relatively modified text, and $I^{t}$ denotes the ground truth target image. For the given inputs of $I^{r}$ and $T^{r}$, the typical pipeline of an SL-CMR model involves mining what content should be modified in the reference image $I^{r}$ according to the caption $T^{r}$, so as to learn a vison-language compositional embedding that encodes the information required to find the interested target image $T^{r}$. Thus, the main challenges faced by such approaches lie in addressing two issues: ”Where to see?”, which refers to attending to the content in the reference image that needs change, and ”How to change?”, which aims to modify the reference image based on the textual information while preserving the remaining information. In recent years, research on SL-CMR has primarily focused on three aspects: (1) data augmentation, focusing on alleviating the difficulty of triplet labeling and the defect of labeling noise; (2) model architecture, focusing on designing a better vision-language combiner via cross-modal feature alignment and fusion strategies, as well as the design of other novel frameworks that can be plugged; (3) loss optimization, focusing on the design of more reasonable feature combination constraints. Although supervised training relying on these carefully labeled data often offers high performance, SL-CMR inherently faces two shortcomings: 1) annotating such triplets is both difficult and labor-intensive, and 2) the supervised approaches trained on the collected limited and specific triplets are also hard for generalization.
@@
Recently, composed multi-modal retrieval based on zero-shot learning has been proposed to address the above limitations. During the training process, the model is trained solely on easily obtainable image-text pairs, without the need for annotated triplets. Its training process usually revolves around learning the modality transformers, which simulate the combination of visual and linguistic information in testing. As a result, the training and testing phases typically involve different network structures. Thus, a main challenge of ZSL-CMR lies in designing transformation frameworks that are equivalent to supervised learning in the absence of supervision signals, aiming to maximize the zero-shot generalization ability as much as possible. To address this challenge, the academic community has developed strategies across three key aspects: (1) Image-side transformation: this approach focuses on learning the implicit or explicit visual-to-linguistic transformation using images as input. During testing, it converts the reference image into a query that can be integrated with the relative textual information; (2) Text-side transformation: in this approach, text is used as input to simulate image features, constructing a training framework that relies solely on language. During testing, the model directly takes image inputs; (3) External knowledge assistance: this approach explores additional knowledge to enhance details, such as attributes and colors, thereby improving retrieval performance. Despite these efforts, there is currently no comprehensive review that systematically organizes the existing research, challenges, and applications of this emerging field.
@@
Although zero-shot combined multimodal retrieval does not rely on labeled data, its performance is often lower than supervised training, which brings obstacles to the application of the model. To alleviate this problem, some works have proposed a semi-supervised combined multimodal learning paradigm based on automatically generated triple data. In this setting, relying on the relatively easy-to-obtain image-text data, existing SSL-CMR work mainly generates triplet data from two aspects: (1) generating images, such as editing the input reference image according to the conditional text to create the target image; (2) generating text, such as describing the difference caption between the two input images. By generating these triplet data, the model can not only capture the combination of vision and language more accurately during learning, but also avoid the limitations of cumbersome annotation. Although the generated data may be noisy, it also combines the advantages of supervision and zero-shot learning, which is a promising direction.
Despite the rapid advancements in composed multi-modal retrieval, numerous challenges remain to be addressed. Issues such as fine-grained vision-language composition, retrieval efficiency in large-scale datasets, and generalization to unseen data require further exploration. Moreover, as real-world applications continue to expand, integrating CMR with cutting-edge technologies such as foundation models and generative AI holds great promise for enhancing retrieval performance. In this paper, we provide a comprehensive review of the existing research in CMR, summarize key challenges, and discuss potential future directions. We hope this work serves as a valuable reference for researchers and practitioners in the field, fostering further innovation in next-generation multi-modal retrieval systems.
@@ 2 Overview
In this section, we provide a concise overview of the taxonomy of current Composed Multi-modal Retrieval (CMR) methods. To better track the latest advancements in CMR, this survey categorizes existing methods from both technical and application perspectives. (1) From a Technical Perspective: we summarize different paradigms of multimodal compositional learning, including sophisticated module designs and innovative solutions. It covers the state-of-the-art techniques that address the challenges of understanding and integrating information from multiple modalities. (2) From an Application Perspective: we examine various application scenarios where multimodal compositional retrieval is employed, such as fashion e-commerce, video platforms, geospatial data, and others. It highlights how these methods are applied in real-world contexts, showcasing their versatility and practical utility. By providing a comprehensive summary and reference for researchers in both the technical and application aspects, this survey aims to facilitate deeper thinking of this field and promote further development in multimodal compositional retrieval. The insights gained from this survey can guide future research and innovation, helping to overcome the inherent complexities and challenges in this rapidly evolving area.
From a technical perspective, current CMR methods can be grouped into three main paradigms: supervised learning, zero-shot learning, and semi-supervised learning. Each paradigm addresses the integration of multiple modalities for effective retrieval, with distinct strategies and trade-offs.
Supervised Learning (SL-CMR) relies on annotated triplet data, such as (reference image $V_{ref}$, modification text $T_{mod}$, target image $V_{tar}$), to learn how to combine input modalities so that the combined content is semantically aligned with the target image. This process can be summarized as:
$X_{combined}=Combiner(V_{ref},T_{mod}),$ (1)
@@
where $Combiner(\cdot)$ denotes the combination of visual and textual inputs, and $Supervisor(\cdot)$ enforces semantic alignment between the combined content and the visual target Supervised learning benefits from explicit supervision, enabling precise modality combination and high performance. However, it requires large-scale, labor-intensive annotation, which can limit its generalization to new domains.
Zero-Shot Learning (ZSL-CMR) overcomes the need for annotated triplets by using readily available image-text pairs to simulate modality combinations. The model learns to transform one modality into another, e.g., mapping a reference image to a textual representation:
$T_{ref}^{*}=Converter(V_{ref}),$ (2)
and then combines this with the modification text to form a composite query. This shifts the task to cross-modal text-to-image retrieval. While ZSL-CMR offers easy data acquisition and better generalization, it lacks direct supervision for modality combination, leading to lower performance on specific tasks compared to supervised learning.
Semi-Supervised Learning (SSL-CMR) combines elements of both supervised and zero-shot approaches. It uses modality generation (e.g., image-to-text or text-to-image) to automatically create triplet data from easily accessible image-text pairs $(V,T)$:
$\hat{V}_{ref},\hat{T}_{mod},\hat{V}_{tar}=Generator(V,T).$ (3)
The model then learns to combine visual and linguistic information, accounting for any noise in the generated data. Semi-supervised learning achieves high performance through automated data construction but faces challenges in generating large volumes of high-quality triplets, especially those with subtle semantic differences. Existing work highlights that generated content often contains errors and noise, which remains an ongoing issue.
Next, we will introduce multimodal combined retrieval based on supervised, zero-shot, and semi-supervised learning in Sections 3, 4, and 5, respectively. Then, in Section 6, we will introduce the applications of combined multimodal retrieval in different scenarios, including related datasets, evaluation metrics, and method performance. Finally, in Section 7, we conclude the paper and look forward to future research directions.

### v2

@@ Abstract
The burgeoning volume of multi-modal data necessitates advanced retrieval paradigms beyond unimodal and cross-modal approaches. Composed Multi-modal Retrieval (CMR) emerges as a pivotal next-generation technology, enabling users to query images or videos by integrating a reference visual input with textual modifications, thereby achieving unprecedented flexibility and precision. This paper provides a comprehensive survey of CMR, covering its fundamental challenges, technical advancements, and applications. CMR is categorized into supervised, zero-shot, and semi-supervised learning paradigms. We discuss key research directions, including data construction, model architecture, and loss optimization in supervised CMR, as well as transformation frameworks and linear integration in zero-shot CMR, and semi-supervised CMR that leverages generated pseudo-triplets while addressing data noise/uncertainty. Additionally, we extensively survey the diverse application landscape of CMR, highlighting its transformative potential in e-commerce, social media, search engines, public security, etc. Seven high impact application scenarios are explored in detail with benchmark data sets and performance analysis. Finally, we further provide new potential research directions with the hope of inspiring exploration in other yet-to-be-explored fields. A curated list of works is available at: Awesome Composed Multi-modal Retrieval.
@@
Composed Multi-modal Retrieval, Vision-Language Semantic Alignment, Multi-modal Semantic Combination
@@ 1 Introduction
Nowadays, the unprecedented prosperity of social media, short video platforms, and e-commerce (such as Facebook, Weibo, YouTube, Amazon, and Alibaba) has greatly promoted the rapid growth of multi-modal data, encompassing various modalities such as texts, images, and videos. Faced with this data flood composed of heterogeneous information, content-based retrieval, as a key technology to search and utilize these vast resources [1], not only plays a core role in the field of e-commerce, such as achieving accurate matching of goods, but also greatly enriches the user experience of the social media, allowing users to quickly and accurately find the content of interest. Due to its widespread applications in everyday life that almost impact everyone, content-based retrieval has attracted great attention from both academia and industry [2, 3].
@@
Generally, the evolution of content-based retrieval technology has witnessed the transformation from Unimodal Retrieval (UR) to Cross-modal Retrieval (CR), and then to Composed Multi-modal Retrieval (CMR). Compared with early-stage unimodal retrieval, which was limited to querying information within the same modality [4, 5, 6], as shown in Fig. 1(a1), cross-modal retrieval has achieved remarkable accuracy and widespread application in the present era. This enables the search for semantically relevant content in one modality based on the instance query from another modality [7, 8, 9, 10, 11], e.g., using text search on images in Fig. 1(a2), allowing users to make full use of these heterogeneous data. In recent years, composed multi-modal retrieval has emerged as a thriving content-based retrieval technology. Within this technical framework [12], as depicted in Fig. 1(a3), the system aims to discover images/videos that not only bear resemblance to the given reference image/video but also allow for specific modifications based on the provided textual feedback from the user. In this sense, CMR pioneers an advanced level of interactive and conditional retrieval mechanisms, leveraging deep integration of visual and linguistic information. This integration greatly enhances the flexibility and precision of user-expressed search intents, injecting new vitality into domains such as internet search and e-commerce. Consequently, CMR exhibits vast potential and far-reaching impact as the next-generation content-based retrieval engine in real-world application scenarios.
Figure 1: (a) The evolution of content-based retrieval technology. (b) In the current research on composed multimodal retrieval (CMR), three main paradigms have been developed. (c) The CMR applications, broadly categorized based on application scenarios and image domains.
@@
A core of CMR is that it requires a synergistic understanding and composition of both input vision and language information as the multi-modal query. The earliest closely related studies of CMR are in the field of attribute-based fashion image retrieval [13, 14, 15, 16, 17, 18, 19], where the key difference is that the textual feedback in attribute-based fashion image retrieval is limited to the predefined attribute value (e.g., ‘mini’, ‘white’, ‘red’), while CMR is the natural language with multiple words (e.g., ‘showing this animal of the input image facing the camera under sunlight’), which is more flexible yet challenging. The pioneering CMR works are proposed in [20], where the input query is specified in the form of an image plus some natural language that describes desired modifications to the input image, leading to a series of subsequent approaches. Current research in CMR is primarily focused on three paradigms: (1) supervised learning-based CMR (SL-CMR), which focuses on how to design a better combination mechanism of vision and language through supervised training of annotated data; (2) zero-shot learning-based CMR (ZSL-CMR), which focuses on how to simulate and build a visual-linguistic multi-modal information combination framework without annotated data; and (3) semi-supervised learning-based CMR (SSL-CMR), which focuses on how to enhance the learning of visual-linguistic combination via generated pseudo-labeling data.
@@
For the SL-CMR pipeline, a notable characteristic is the requirement of annotated triplet data $(I_{r},T_{m},I_{t})$, which denotes the reference query image, the modified text, and the ground-truth target image, respectively. As illustrated in Fig. 1(b1), for the given inputs $I_{r}$ and $T_{m}$, SL-CMR involves mining the content that should be modified in the reference image $I_{r}$ according to the text $T_{m}$, so as to learn a multi-modal compositional embedding to find the interested target image $I_{t}$. Thus, the challenges faced by SL-CMR mainly lies in addressing two issues: “Where to see”, which refers to attending to the content in the reference image that needs change, and “How to change”, which aims to modify the reference image based on the textual information while preserving the remaining information. In recent years, research on SL-CMR has primarily focused on three aspects: (1) data construction [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], focusing on labeling triples with accurate semantic difference descriptions; (2) model architecture [20, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 21, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65], focusing on designing a better vision-language combiner via cross-modal feature alignment and fusion strategies, as well as the design of other novel frameworks that can be plugged; (3) loss optimization [20, 66, 67, 21, 40, 68, 69, 48, 62, 70, 71, 72, 73, 44], focusing on the design of more reasonable feature combination constraints. Although supervised training relying on these carefully labeled data often offers high performance, SL-CMR inherently faces two shortcomings: 1) annotating such triplets is both difficult and labor-intensive, and 2) the supervised approaches trained on the collected limited and specific triplets are also hard for generalization.
@@
Recently, ZSL-CMR has been proposed to address the above limitations, where the model is trained solely on easily obtainable large-scale image-caption pairs or unlabeled images. As depicted in Fig. 1(b2), its training process usually revolves around learning the modality converters that simulate the combination of visual and linguistic information in the test. The training and testing phases typically involve different network structures. Thus, a main challenge of ZSL-CMR lies in designing transformation frameworks that achieve accurate vision-language combination in the absence of supervision signals, aiming to maximize the zero-shot generalization ability. To address this challenge, the academic community has developed strategies across three key aspects: (1) image-side transformation [74, 75, 76, 23, 12, 77, 78, 79, 80, 81]: this approach focuses on learning the implicit or explicit visual-to-linguistic transformation using images as input. During testing, it converts the reference image into a query that can be integrated with the relative textual information; (2) text-side transformation [82, 83, 84, 85, 86]: in this approach, text is used as input to simulate image features, constructing a training framework that relies solely on language. During testing, the model directly takes image inputs; (3) linear interpolation [87, 88, 89, 90]: this approach explores the simple yet effective linear weighted combination strategies of visual and textual features.
@@
Although zero-shot approaches do not rely on labeled data, their performance is often lower than supervised training, which brings obstacles to the application of the model. To alleviate this problem, as shown in Fig. 1(b3), some CMR works have proposed the semi-supervised learning paradigm based on generated pseudo-labeling data. In this setting, relying on the relatively easy-to-obtain image-text data, existing SSL-CMR works mainly generate triplet data from two aspects: (1) generating text [91, 92, 71, 93, 94, 95, 96], such as describing the difference caption between the two input images; and (2) generating images [97, 98, 73], such as editing the input reference image according to the conditional text to create the target image. Besides, alleviating the noise in generated data is also a focus of this paradigm [91, 94, 98, 36, 92, 96, 72, 99, 100, 101, 23, 32]. In this way, the model can not only capture the combination of vision and language more accurately during learning, but also avoid the limitations of cumbersome annotation. Although the generated data may introduce uncertainty, this paradigm combines the advantages of supervision and zero-shot learning, which is a promising direction.
Research in CMR has vast application potential. As illustrated in Fig. 1(c), it can be broadly categorized based on application scenarios and image domains, including fashion and E-commerce images [26, 27, 28, 32, 102, 103, 104, 105, 106], natural images [20, 21, 23, 96, 91, 98, 107], videos [29, 108, 84, 74, 30], remote sensing images [109, 110], person images [111], sketch images [112], and interactive conversation [28, 113, 114, 115, 116, 117, 118, 119, 120, 121]. The specific application can be personalized product shopping, media search, event discovery, environmental monitoring, law enforcement, customer service bots, and so on. In summary, CMR represents a paradigm shift in search systems by integrating visual and textual modalities. These systems enable fine-grained, context-aware, and user-centric searches across diverse domains, offering significant improvements in both retrieval accuracy and user satisfaction.
Compared to related surveys [122, 123] that appeared around the same period of early 2025 on arXiv, the novelty and contributions of this survey are: (1) We comprehensively summarize methods and techniques of visual-language retrieval in recent years, especially cross-modal semantic alignment and combination learning, covering more than 250 works, providing a more in-depth summary for the community. In particular, for the first time, we comprehensively review 6 types of combiners, introduce more than 7 insighful losses, 3 zero-shot combinations, including a new linear interpolation perspective, noise and uncertainty considerations in semi-supervision; (2) We comprehensively emphasize 7 applications and potential scenarios of CMR, and provide more than 26 commonly used datasets and detailed performance statistics for each application, which is the most covered so far; (3) We further provide new potential research directions and guidance for CMR, hoping to inspire exploration in other unexplored scenarios.
@@ 2 Overview
In this section, we provide a concise overview of the taxonomy of current Composed Multi-modal Retrieval (CMR) methods. To better track the latest advancements in CMR, this survey categorizes existing methods from both technical and application perspectives. (1) From a Technical Perspective: we summarize different paradigms of multimodal compositional learning, including sophisticated module designs and the state-of-the-art innovative solutions. (2) From an Application Perspective: we examine various application scenarios where CMR is employed, such as fashion e-commerce, video platforms, geospatial data, and others. It highlights how these methods are applied in real-world contexts, showcasing their versatility and practical utility. By providing a comprehensive summary and reference for researchers in both the technical and application aspects, this survey aims to facilitate deeper thinking in this field and promote further development.
2.1 CMR Methodology
From the technical perspective, current CMR methods can be grouped into three main paradigms: supervised learning (SL), zero-shot learning (ZSL), and semi-supervised learning (SSL). Each paradigm addresses the understanding and integration of multiple modalities for effective retrieval, with distinct strategies and trade-offs.
SL-CMR relies on annotated triplet data, i.e., (reference image $I_{r}$, modification text $T_{m}$, target image $I_{t}$), to learn how to combine input modalities so that the composed content is semantically aligned with the target image. This process can be summarized as:
$X_{composed}=Combiner(I_{r},T_{m}),$ (1)
@@ 2.1 CMR Methodology
where $Combiner(\cdot)$ denotes the combination of visual and textual inputs, and $Optimizer(\cdot)$ enforces semantic alignment between the composed content and the visual target. This paradigm benefits from explicit supervision, enabling precise modality combination and high performance. However, it requires large-scale, labor-intensive annotations, which limit its generalization to new domains.
ZSL-CMR overcomes the need for annotated triplets by using readily available image-text pairs. Here, they mainly focus on converting one modality into another, e.g., mapping a reference image to a textual representation:
$T_{r}^{*}=Converter(I_{r}),$ (2)
and then combines this with the modification text $T_{m}$ to form a composite query. This shifts the task to cross-modal text-to-image retrieval. While ZSL-CMR offers easy data acquisition and better generalization, it lacks direct supervision for modality combination, leading to lower performance on specific tasks compared to supervised learning.
SSL-CMR combines elements of both supervised and zero-shot approaches. It uses modality generation (e.g., image-to-text or text-to-image) to automatically create triplet data from easily accessible image-text pairs $(I,T)$:
$\hat{I}_{r},\hat{T}_{m},\hat{I}_{t}=Generator(I,T).$ (3)
Although this paradigm yields high performance, it faces challenges in generating large volumes of high-quality triplets, especially those with subtle semantic differences. Existing work highlights that generated content often contains errors and noise, which remains an ongoing issue.
Figure 2: Composed Multi-modal Retrieval Methods Taxonomy.
2.2 CMR Applications
CMR can be broadly categorized based on application scenarios and image domain differences. Specifically, (1) In fashion and e-commerce fields, CMR enables users to search for products by combining reference images with descriptive text, improving personalization and shopping accuracy. (2) In natural image domains, CMR helps users refine searches with contextual cues (e.g., ”same scene in autumn”), supporting creative and personal content discovery. (3) In composed video retrieval, composed queries allow users to locate dynamic content with greater precision by specifying conditions like time or activity. (4) In remote sensing image domains, CMR deals with satellite imagery, where text adds geographic or temporal specificity for tasks like urban planning or disaster monitoring. (5) Composed person retrieval combines visual input from a specific person with attribute-based descriptions to support identity matching in surveillance or public safety. (6) Sketch-based CMR leverages user-drawn inputs and textual hints, particularly useful in design and creative industries, or law enforcement when photos are unavailable. (7) Finally, in the interactive conversational scenarios, CMR allows users to refine queries through multi-turn dialogue, enabling more flexible and accurate searches across various domains.
Next, we will introduce composed multi-modal retrieval based on supervised, zero-shot, and semi-supervised learning in Sect. 3, 4, and 5, respectively. Then, in Sect. 6, we will introduce the CMR applications in different scenarios, including related benchmark datasets, evaluation metrics, and method performance. Finally, in Sect. 7 and Sect. 8, we look forward to future research directions and conclude the paper, respectively.